# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 02 Mar 2026 v1

@@ B.2 Experiment B: Approximate Exchangeability and Variance Inflation
Setup.
To model mild departures from ideal exchangeability (e.g., weak dependence, contamination, or partial reuse of samples), we inflate the intrinsic covariance by a factor $(1+\varepsilon)$ in the linearized surrogate:
$\mathrm{Cov}(\xi_{t})=\frac{1}{b}\,(1+\varepsilon)\,M^{\star},\qquad\varepsilon\in[0,1].$
This captures the practical effect of “extra correlation” as an increase in effective diffusion amplitude.
Benchmark.
Linearity of the Lyapunov operator implies
$\Sigma(\varepsilon)=(1+\varepsilon)\Sigma(0),$
so every stationary marginal variance inflates linearly in $\varepsilon$.
What the figure shows.
Figure 10 confirms the predicted linear inflation: the entire stationary covariance scales as though the system had a higher effective temperature $\tau_{\mathrm{eff}}=(1+\varepsilon)\eta/b$. This provides a compact operational summary of small exchangeability violations: to first order they act like “heating” the Fisher–Lyapunov equilibrium.

### 03 Apr 2026 v2

@@ B.2 Experiment B: Approximate Exchangeability and Variance Inflation
To model mild departures from ideal exchangeability (e.g., weak dependence, contamination, or partial reuse of samples), we inflate the intrinsic covariance by a factor $(1+\varepsilon)$ in the linearized surrogate: $\mathrm{Cov}(\xi_{t})=(1/b)(1+\varepsilon)\,G^{*}(\theta^{\star})$, $\varepsilon\in[0,1]$. This captures “extra correlation” as increased effective diffusion amplitude. Linearity of the Lyapunov operator then implies $\Sigma(\varepsilon)=(1+\varepsilon)\Sigma(0)$, so every stationary marginal variance inflates linearly in $\varepsilon$. Figure 10 confirms this prediction: the entire stationary covariance scales as though the system had a higher effective temperature $\tau_{\mathrm{eff}}=(1+\varepsilon)\eta/b$. To first order, small exchangeability violations act like “heating” the Fisher–Lyapunov equilibrium.
@@
Appendix C Reproducibility and Code Availability
All experiments are reproducible from the accompanying code archive. The Python suite (Python_Codes/) includes a master script run_all.py that executes all experiments sequentially. Every script uses explicit random seeds (global seed =7=7) for deterministic reproducibility; small platform-level floating-point differences may produce minor variations ($\lesssim 1\%$) in Monte Carlo estimates.
Software environment.
Python $\geq$3.9; NumPy, SciPy, Matplotlib, Pandas. Tested under Python 3.10.12, NumPy 2.2, SciPy 1.15, Matplotlib 3.10. Total wall-clock time: $\approx$4 minutes on a single CPU core.
Script inventory.
Table 6 maps each script to the figures and tables it generates.
Table 6: Code inventory: scripts, outputs, and paper cross-references.
Script Experiment Figures / Tables Sec.
partA_batch_scaling.py 1 (Batch scaling) Fig. 2 8.1
partB_anisotropic_alignment.py 2 (Angle sweep) Fig. 3 8.1
partC_rotation_sweep.py 3 (Rotation) Figs. 57 8.3
bounds.py 4–6 (OU plateau, rate, directional) Tab. 4, Figs. 48, Tab. 5 8.18.3
partD_misspecification.py A (Misspecification) Fig. 9 App. B
partE_eps_exchangeable.py B ($\varepsilon$-exchangeable) Fig. 10 App. B
The script run_all.py executes all of the above in sequence. The script partF_optimal_batching.py produces additional exploratory figures not included in the paper.
Shared parameters ($d=10$ experiments).
Hessian eigenvalues log-spaced in $[1,100]$; $F^{\star}=H^{\star}$ (quadratic Gaussian model). Anisotropic gradient covariance $G^{\star}$ drawn as a random SPD matrix with eigenvalues in $[0.5,5]$ (seed =7=7); isotropic baseline $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Constant-step experiments use $\eta=0.5/\lambda_{\max}(H)$, $b=64$, $\tau=\eta/b$; decaying-step experiments use $\eta_{t}=1.2/(t+50)$, $b=64$.
Verification.
All scripts were executed to completion with zero errors. Numerical outputs confirm: (i) log-log slopes $\approx-1$ for the $1/N$ risk decay; (ii) OU stationary variances match Lyapunov predictions to $1\%$; (iii) directional variance ratios align with the eigenstructure of $G^{\star}$; (iv) misspecification experiments show graceful degradation to the sandwich/Godambe prediction.